# Functional classification of metabolic networks

### v1

@@ Conclusions
Recent experimental and computational developments provide us with the opportunity to develop functional classifications of chemical reaction networks grounded in physical principles. Here, we introduce a framework for the classification of these networks using differences between nullspaces of their stoichiometric matrices. In the human gut, this framework enables us to discover metabolic processes that describe groups of bacteria with similar steady-state fluxes and conservation laws. The generality of this framework, from chemical reaction networks in bacteria to human tissues to planetary atmospheres, can lead to the development of a universal atlas of chemical reaction networks in which systems across length scales must reside.
Acknowledgements
We are grateful to Gene-Wei Li and Leonid Mirny for helpful discussions. This work was supported by a MathWorks Science Fellowship (J.R.), the National Science Foundation Graduate Research Fellowship Program under Grant No. 2141064 (J.R.), the MathWorks Professorship Fund (J.D.), Alfred P. Sloan Foundation Grant G-2021-16758 (J.D.), and through Schmidt Sciences LLC (Polymath award to J.D.). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Methods
Without proper conditioning, any chemical reaction network may not be physically admissible. We ensure all networks considered here satisfy flux and stoichiometric consistency using methods available in the COBRA Toolbox [84, 85, 86, 82, 87] with Gurobi v11.0.2 [83]. Flux consistency implies that each process of a chemical reaction network is active, realizing nonzero flux, in at least one flux distribution [84, 85, 87]. We remove flux inconsistent internal (non-exchange) processes to ensure we only consider active metabolic processes [85, 86]. Similarly, reaction databases may contain stoichiometric inconsistencies [88], where the stoichiometry of processes is inconsistent with conservation of mass. To that end we identify and correct stoichiometric inconsistent internal processes [86].
@@
We compute silhouette scores as goodness of clustering scores using the Julia Clustering.jl package [90, 35]. We identify clusters for metabolic distance matrix by hierarchical clustering with Ward linkage (Minimum Increase of Sum of Squares) and obtain clusters with a tree-cut to the desired level of granularity. To assign identities to clusters we perform a cluster comparison analysis for each cluster by masking assignments to inclusion within a cluster of interest and identifying which processes(s) are enriched or depleted in the cluster. Since certain processes are always jointly present in the networks, we group processes if they co-occur in the same manner across all our networks. We identify those processes that minimize the variation of information [59] between inclusion masked assignments and process presence.
@@
$I(X,Y)$ $=\sum_{k,k^{\prime}}P(k,k^{\prime})\log\frac{P(k,k^{\prime})}{P(k)P(k^{\prime})},$
$P(k,k^{\prime})$ =#$=\frac{\#|X_{k}\cap Y_{k^{\prime}}|}{\#|S|}.$
@@ I The square root of a metric is a metric
    $\tilde{d}(x,y)$ $=\sqrt{d(x,z)}\leq\sqrt{d(x,y)+d(y,z)}\leq\sqrt{d(x,y)+d(y,z)+2\sqrt{d(x,y)d(y,z)}}$
    $=\sqrt{\left(\sqrt{d(x,y)}+\sqrt{d(y,z)}\right)^{2}}=\sqrt{d(x,y)}+\sqrt{d(y,z)}=\tilde{d}(x,y)+\tilde{d}(y,z).$
@@ II Broken conservation laws
Consider the simple three-component ($A,B,C$) closed chemical reaction network [38] with stoichiometric matrix
@@
Re-closing the system by introducing three analogous chemical species in the environment provides the augmented stoichiometric matrix
@@
Figure S1: Dimension of right nullspace colors direction in R-Grassmann embedding. Bacterial network size (A) is linearly predictive of the right nullspace dimension, in contrast to the left nullspace (B). (C) Colorings the sgd-MDS embedding of the R-Grassmann distance matrix by network size and nullspace dimension are consistent with directions corresponding to nullspace size.
@@
Figure S5: Embeddings of distances highlight that metabolic groups are robust to the choice of embedding procedure. (A) Colors correspond to metabolic groups identifies in Fig. 5. t-SNE complexity and learning rate are set to 25 and 50, respectively. UMAP # neighbors and minimum distance are set to 10 and 2, respectively. (B) For each partition based on metabolic group inclusion, we perform a local sgd-MDS embeddings against the 19 nearest partitions based on the presence of metabolic processes. These embeddings demonstrate the local binary partition structure of the identified metabolic groups. Local sgd-MDS parameters: $\varepsilon=0.1$, $t_{s}=10,000$, $\tau_{s}=10,000$, $\delta=10^{-8}$.

### v2

@@ Conclusions
Recent experimental and computational developments provide us with the opportunity to develop functional classifications of chemical reaction networks grounded in physical principles. Here, we introduce a framework for the classification of these networks using differences between nullspaces of their stoichiometric matrices. In the human gut, this framework enables us to discover metabolic processes that describe groups of bacteria with similar steady-state fluxes and conservation laws. The generality of this framework, from chemical reaction networks in bacteria to human tissues to planetary atmospheres, can lead to the development of a universal atlas of chemical reaction networks in which systems across length scales must reside. Moreover, by recasting complex nonlinear dynamical systems as effective transport networks in the form of (1), the above methodology becomes broadly applicable to other dynamical phenomena.
@@ Methods
Without proper conditioning, any chemical reaction network may not be physically admissible. We ensure all networks considered here satisfy flux and stoichiometric consistency using methods available in the COBRA Toolbox [87, 88, 89, 85, 90] with Gurobi v11.0.2 [86]. Flux consistency implies that each process of a chemical reaction network is active, realizing nonzero flux, in at least one flux distribution [87, 88, 90]. We remove flux inconsistent internal (non-exchange) processes to ensure we only consider active metabolic processes [88, 89]. Similarly, reaction databases may contain stoichiometric inconsistencies [91], where the stoichiometry of processes is inconsistent with conservation of mass. To that end we identify and correct stoichiometric inconsistent internal processes [89] (SI Appendix).
@@
We compute silhouette scores as goodness of clustering scores using the Julia Clustering.jl package [93, 38]. We identify clusters for metabolic distance matrix by hierarchical clustering with Ward linkage (Minimum Increase of Sum of Squares) and obtain clusters with a tree-cut to the desired level of granularity. To assign identities to clusters we perform a cluster comparison analysis for each cluster by masking assignments to inclusion within a cluster of interest and identifying which processes(s) are enriched or depleted in the cluster. Since certain processes are always jointly present in the networks, we group processes if they co-occur in the same manner across all our networks. We identify those processes that minimize the variation of information [62] between the inclusion masked assignments and process presence.
@@
$I(X,Y)=\sum_{k,k^{\prime}}P(k,k^{\prime})\log\frac{P(k,k^{\prime})}{P(k)P(k^{\prime})},\ P(k,k^{\prime})=\frac{\#|X_{k}\cap Y_{k^{\prime}}|}{\#|S|}.$
@@
Data Availability
All study data and relevant codes are available on a public Github repository at https://github.com/jrysyrj.
Acknowledgments
We are grateful to Gene-Wei Li and Leonid Mirny for helpful discussions. This work was supported by a MathWorks Science Fellowship (J.R.), the National Science Foundation Graduate Research Fellowship Program under Grant No. 2141064 (J.R.), the National Science Foundation DMR/MPS-2214021 (J.D.), the MathWorks Professorship Fund (J.D.), Alfred P. Sloan Foundation Grant G-2021-16758 (J.D.), and through Schmidt Sciences LLC (Polymath award to J.D.). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ I The square root of a metric is a metric
    $\tilde{d}(x,y)$ $=\sqrt{d(x,z)}\leq\sqrt{d(x,y)+d(y,z)}$
    $\leq\sqrt{d(x,y)+d(y,z)+2\sqrt{d(x,y)d(y,z)}}$
    $=\sqrt{\left(\sqrt{d(x,y)}+\sqrt{d(y,z)}\right)^{2}}$
    $=\sqrt{d(x,y)}+\sqrt{d(y,z)}$
    $=\tilde{d}(x,y)+\tilde{d}(y,z).$
@@
Note that this implies that any even root of a metric is also metric.
@@ II Broken conservation laws
Here we recapitulate the results of the simple three-component closed chemical reaction network reported in [41]. The corresponding stoichiometric matrix is
@@
Re-closing the system by introducing three analogous components in the environment provides the augmented stoichiometric matrix
@@
Figure S1: Dimension of right nullspace colors direction in R-Grassmann embedding. Bacterial network size (A) is linearly predictive of the right nullspace dimension, in contrast to the left nullspace (B). (C) Colorings the sgd-MDS embedding of the Grassmann distance matrix by network size and nullspace dimension are consistent with directions corresponding to nullspace size.
@@
Figure S5: Embeddings of distances highlight that metabolic groups are robust to the choice of embedding procedure. (A) Colors correspond to metabolic groups identifies in Fig. 5. t-SNE complexity and learning rate are set to 25 and 50, respectively. UMAP # neighbors and minimum distance are set to 10 and 2, respectively. (B) For each partition based on metabolic group inclusion, we compute a histogram of variation of information (VI) distances to assess the uniqueness of label assignments. Bins corresponding to metabolic processes that achieve a minimal VI against the partitions are identified with a four-point star. We observe that some, but not all, metabolic processes correspond to small distances with low frequency.